A man who can swim at the rate of 2km/hr, crosses a river to an exactly opposite point on the other bank by swimming in a direction of 120 degree to the flow of the water in the river. The velocity of the water current in km/hr is __ ??

Grade:11

2 Answers

raj mishra
11 years ago

ans 1km/hr

let v be the velocity of water

then v= Vm* cos 60

2*1/2=1 km/h

shivang belwariar
11 years ago

look

he can swim will always mean velocity of man relative to river .

Vmr = Vm - Vr

 

taking components -

Vmrcos120 = Vmcos90 - Vrcos0

 

2*(-1/2) = -Vr

 

SO , Vr = 1 km/hr

 

plz like my answer